Shift Summary 8/28/99 Evening shift

Jianbo Wang (jwang@physics.Arizona.EDU)
Sat, 28 Aug 1999 21:58:25 -0700 (MST)

Shift summary for Evening shift, 8/28/99

On Shift: Theo, Jianbo
Sydney was here for the whole shift.
Bob T was here for pipeline power supply.
Nick came over here after his MINOS meeting and was working on his KAON99
paper continuously.

Hours of beam available at all intensities: 8.25h
Hours of beam available at nominal intensity:8.25h
Hours of E832 data (special runs not included): 3.5h
Hours of special runs (laser scans, muon runs...):0h

Accounting of lost time: dc2-VME problem, swap a bad dc2.
Access for swaping a bad pipeline power supply.

========================================================================
Runs ended this shift:
14368 BEAM_832 - 5.7M events

Tapes brought to FCC this shift:

Accesses during shift: Replace a bad pipeline power supply.

========================================================================
Overview of shift:

Shift from hell, many problems-- dc2-VME, pipeline power supply, DDD veto

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14366 15:30 in progress (later overwritten, see below)

16:20 plane 4 dies caused by stream 4 problems, paged
Sydney,and JasonL.

16:55 Vivian swaps the remaining two memory
modules (the first module was swapped
earlier) between planes 2 and 4 for stream 4,
and replaces DC2 for plane 4 stream 4.

14368 17:10 Start the run, overwrote run 14366 (35 spills)
QK*345-347 QKB120 after an attempt to write EOF
markers failed.

60 18:30 pipeline crate 7 is full. Initialize this crate.
Kfrend. Then crate 3 is full. Initialize
all crates but crate 6 is off. Check the status of
crate 6, we found that voltage interlock is
dropped.

18:45 Paged BobT

19:15 BobT and Sydney made a access.

Found power supply for pipeline crate #6 was bad.
Swapped and put in spare. Also Sydney fixed
Boot scripts for downstream x_term near CsI. We
can now run pipe.tcl there.

21:05 BobT and Sydney out of access.

185 21:15 Beam returns.

210 21:49 Ktev2 DDD is busy -- gateway process died.

215 21:56 Stop the run 14368. Drt_fresh_start.

21:57 Onsplit request tape QKMS15 at tps3d2 on ktev1

22:05 Start run 14369 -- gateway process on ktev2 dies.

From murmur message, it turns out. problem
W/ PII driver check caught error
reading from first memory module.

Paged Viv. She directs us to
reseat plane 2 stream 4 memory modules,
drt_fresh_start.

22:45 Start run 14370, and the problem is fixed.

23:30 Onsplit request tape QKNL28 at tps3d2 at ktev2.

========================================================================
Follow up Notes on Previous Shift:
o kept our eyes on channel 450.
It was reasonably stable during the whole shift.
========================================================================
Items to Follow up on Next Shift:

o After changing the tapes, please do muon run.

========================================================================